Consider a pair of discrete random variables X and Y. suppose that the marginal distribution of X is given by the table below. x 0.20 0.80 Suppose furthermore that the conditional distributions of tables below... given X are given by the two y0.20 0.80 0.60 0.40 Enter the joint probability mass function of X and Y into the table below .r
Enter the joint probability mass function of X and Y into the table below. Check
Answer #1
x
1 2
y 1 0.04 0.16
2 0.48 0.32

formula

P(X=x, Y =y) = P(Y=y|X= x) P(X = x)

x
1 2
y 1 =0.2*0.2 =0.8*0.2
2 =0.8*0.6 =0.8*0.4

Please rate